OpenStudy (anonymous):

What is the solution to the equation log4 x – log4 2 = 3 ?

OpenStudy (lalaly):

\[\log_4(\frac{x}{2})=3\]\[4^3=\frac{x}{2}\]solve for x

OpenStudy (lalaly):

\[\log_b A-\log_bC= \log_b \frac{A}{C}\]

OpenStudy (anonymous):

log4 x – log4 2 = log4 (x/2) so: log4 (x/2) = 3 and 4^3 = (x/2)

OpenStudy (anonymous):

x=128

OpenStudy (anonymous):

yup